OpenStudy (anonymous):

Help me Please 1. Find the inverse of f(x) = √(x2 - 3). Make sure to show your work. 2. Find the inverse of f(x) = 7x2 + 12. Make sure to show your work. 3. Explain how to determine if a given relation is a one to one function. You may use examples to help in your explanation.

OpenStudy (anonymous):

@Squirrels

OpenStudy (bibby):

replace f(x) with y switch x and y solve for y

OpenStudy (bibby):

here \( f(x) = \sqrt{(x^2 - 3)}\) \(y = \sqrt{(x^2 - 3)}\) \(x = \sqrt{(y^2 - 3)}\) what would our first step be?

OpenStudy (anonymous):

so square each side to get rid of the square root

OpenStudy (bibby):

noice, noice so now we have \(x^2 = y^2 - 3\)

OpenStudy (anonymous):

yes

OpenStudy (bibby):

so, what would our next step be?

OpenStudy (anonymous):

-3 each side?

OpenStudy (anonymous):